OpenStudy (anonymous):

How can you add eight 8's to get the number 1,000?

OpenStudy (agent0smith):

If you add 8 125 times, you'll get to 1000. 8x125 = 1000

OpenStudy (anonymous):

There are only 8 8's.. you cannot add them 125 times..

zepdrix (zepdrix):

I'm not sure I understand the question correctly, but did you mean, perform 8 operations with the number 8 and reach 1000? Is this the idea? \[\large (8\cdot8)+(8\cdot8)=128\] \[\large 128 \cdot 8 = 1024\]\[\large 1024-8-8-8=1000\]

OpenStudy (agent0smith):

Ah, yes i mis-read that.

OpenStudy (anonymous):

888 + 88 + 8 + 8 + 8 = 1000 :D

OpenStudy (agent0smith):

Just read it as "how can you add 8's to get to 1000"

OpenStudy (anonymous):